OpenStudy (anonymous):

what is the eccentricity of (x^2/9)-(y^2/4)=1?

OpenStudy (anonymous):

well is it x to the 2 9ths or it it x to the 2 divided by 9

OpenStudy (anonymous):

\[\frac{ x^2 }{ 9 }-\frac{ y^2 }{ 4 }=1\] that is the equation

OpenStudy (anonymous):

wait I think this is past my level of understanding. I'm sorry and I wish I could help. I don't think I know what eccentricity means. If you still want I could try, if not maybe give the medal to someone else.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

e = c/a

OpenStudy (anonymous):

since it is a hyperbola c^2 = a^2 + b^2

OpenStudy (anonymous):

a = 3 b = 2

OpenStudy (anonymous):

yeah i figure that much but my answer is wrong i wonder what is the right answer?

OpenStudy (anonymous):

c= sqrt (9 +4) = sqrt (13) e = sqrt(13)/3

OpenStudy (anonymous):

well my answer is\[\frac{ \sqrt{5} }{ 3 }\] oh it should \[\sqrt{13}\] not\[\sqrt{5}\] thanks

OpenStudy (anonymous):

If it is an ellipse then ur answer is correct

OpenStudy (anonymous):

But it is Hyperbola

OpenStudy (anonymous):

thanks for clarifying it

OpenStudy (anonymous):

Welcome